OpenStudy (freemap):

Which lines are perpendicular to 5x + 4y = 8? Choose all answers that are correct. A. –5x + 4y = –5 B.5y = –4x – 15 C.5y = 4x + 5 D.–4x + 5y = 10

Vocaloid (vocaloid):

our first step is to convert each equation into y = mx + b form, do you know how to do that?

OpenStudy (freemap):

YES, is one of the answers A?

Vocaloid (vocaloid):

hm, I don't think A is one of the answers, let's walk it through step by step let's start with the original equation 5x + 4y = 8 we can re-arrange this to get 4y = -5x + 8 y = (-5/4)x + 2 the slope of the original equation is (-5/4), so any line that is perpendicular will have a slope of (4/5) check each of the four answer choices to see which one(s) have a slope of 4/5

OpenStudy (freemap):

D

Vocaloid (vocaloid):

D is one of the answers, yes. there's also one more, can you find it? :)

OpenStudy (freemap):

I think so, just a sec

OpenStudy (madhu.mukherjee.946):

@Vocaloid great job

Vocaloid (vocaloid):

thank you ^_^

OpenStudy (dinamix):

its D the answer and i sure about it

Vocaloid (vocaloid):

yup, C and D are both answers

OpenStudy (freemap):

Thanks so much!!!!

OpenStudy (dinamix):

@Vocaloid  @Vocaloid , @freemap